# Simple Semi-supervised Knowledge Distillation from Vision-Language Models via Dual-Head Optimization

### v1

@@ 3.2 Dual-Head Optimization (DHO)
Figure 4: The average cosine similarity and inner product across 10 datasets. $\hat{p}_{\text{CE}}=\hat{p}_{\text{KD}}=\sigma(h(z))$ for SHO.
Gradient conflicts in single-head KD.
One of the single-head KD baselines, logit distillation, described in §3.1, offers a simple and efficient approach for transferring knowledge from VLMs in semi-supervised settings. However, we observe that its performance gain is suboptimal, which we attribute to gradient conflicts between the supervised and KD loss signals—a phenomenon well known to hinder effective feature representations in multi-task learning literature [110, 54, 11]. As illustrated in of Figure 4, the parameter vector $\theta$ of feature extractor $g$ suffers from this issue: the cosine similarity between the gradients often turns negative, i.e., $\mathtt{CosSim}(\nabla_{\theta}\mathcal{L}_{\text{CE}},\nabla_{\theta}\mathcal{L}_{\text{KD}})<0$, indicating misaligned optimization directions.
Gradient analysis of single-head KD.
To understand this, we analyze the gradient w.r.t feature representation $z$. We denote $h(z)=Wz+b$, where $W\in\mathbb{R}^{C\times d}$, $b\in\mathbb{R}^{C}$, and $\hat{p}=\sigma(h(z))$. Then, gradients for a labeled data and their cosine similarity are [^1] Note that this is not an exact cosine similarity, which depends on the full gradients $\nabla_{\theta}\mathcal{L_{\text{CE}}},\nabla_{\theta}\mathcal{L_{\text{KD}}}$ involving the Jacobian $\partial z/\partial\theta$. Also, $(\hat{p}-y)^{\top}(\hat{p}-p)<0\nRightarrow(\hat{p}-y)^{\top}WW^{\top}(\hat{p}-p)<0$. Still, our analysis with observations offers useful intuition on how the discrepancy $(\hat{p}-y)^{\top}(\hat{p}-p)$ can amplify gradient conflicts.:
$\nabla_{z}\mathcal{L}_{\text{CE}}=W^{\top}(\hat{p}-y),\quad\nabla_{z}\mathcal{L}_{\text{KD}}=W^{\top}(\hat{p}-p),$ (4)
$\mathtt{CosSim}(\nabla_{z}\mathcal{L}_{\text{CE}},\nabla_{z}\mathcal{L}_{\text{KD}})=\frac{(\hat{p}-y)^{\top}WW^{\top}(\hat{p}-p)}{||W^{\top}(\hat{p}-y)||_{2}\cdot||W^{\top}(\hat{p}-p)||_{2}}.$ (5)
$y$ denotes the one-hot ground-truth label and $p$ the teacher prediction. In few-shot settings, labeled data is scarce, making supervised signals noisy, while teacher predictions reflect pretraining priors that may not align with the limited labels. This mismatch often leads to a prediction-level conflict, indicated by $(\hat{p}-y)^{\top}(\hat{p}-p)<0$, which rapidly falls below zero during training ( $\times$ in Figure 4). As the dominant eigenvalue of $WW^{\top}$ grows to enhance class separability (Figure 13-(left)), it amplifies the conflict magnitude, causing $\mathtt{CosSim}(\nabla_{z}\mathcal{L}_{\text{CE}},\nabla_{z}\mathcal{L}_{\text{KD}})$ to become strongly negative (Figure 13-(right)). Eventually, this drives $\mathtt{CosSim}(\nabla_{\theta}\mathcal{L}_{\text{CE}},\nabla_{\theta}\mathcal{L}_{\text{KD}})$ down to $-0.6$ ( ).
@@
To mitigate this issue, we propose to decouple the supervised and KD objectives via two independent prediction heads: $h_{\text{CE}}(z)=W_{\text{CE}}z+b_{\text{CE}}$ and $h_{\text{KD}}(z)=W_{\text{KD}}z+b_{\text{KD}}$, with $W_{\text{CE}},W_{\text{KD}}\in\mathbb{R}^{C\times d}$ and $b_{\text{CE}},b_{\text{KD}}\in\mathbb{R}^{C}$. The corresponding losses are:
@@
and the final loss is $\lambda\mathcal{L}_{\text{CE}}+(1-\lambda)\mathcal{L}_{\text{KD}}$, following single-head KD. Denoting $\hat{p}_{\text{CE}}=\sigma(h_{\text{CE}}(z))$ and $\hat{p}_{\text{KD}}=\sigma(h_{\text{KD}}(z))$, the gradients for a labeled data and their cosine similarity become:
$\nabla_{z}\mathcal{L}_{\text{CE}}=W_{\text{CE}}^{\top}(\hat{p}_{\text{CE}}-y),\quad\nabla_{z}\mathcal{L}_{\text{KD}}=W_{\text{KD}}^{\top}(\hat{p}_{\text{KD}}-p),$ (8)
@@
$\mathtt{CosSim}(\nabla_{z}\mathcal{L}_{\text{CE}},\nabla_{z}\mathcal{L}_{\text{KD}})=\frac{(\hat{p}_{\text{CE}}-y)^{\top}W_{\text{CE}}W_{\text{KD}}^{\top}(\hat{p}_{\text{KD}}-p)}{||W_{\text{CE}}^{\top}(\hat{p}_{\text{CE}}-y)||_{2}\cdot||W_{\text{KD}}^{\top}(\hat{p}_{\text{KD}}-p)||_{2}}.$ (9)
Crucially, as shown in $\times$ of Figure 4, the prediction discrepancy term $(\hat{p}_{\text{CE}}-y)^{\top}(\hat{p}_{\text{KD}}-p)$ of DHO remains near zero throughout training. This is because the separated heads $h_{\text{CE}}$ and $h_{\text{KD}}$ are learned independently, without shared projections or alignment constraints. This is further supported by the high-dimensional geometry (e.g., $C>100$): random vectors drawn from a simplex or sphere tend to be nearly orthogonal due to the concentration of measure [92], resulting in weak correlation between their directions. Although spectral amplification occurs for $W_{\text{CE}}W_{\text{KD}}^{\top}$ (Figure 13-(left)), the discrepancy term $(\hat{p}_{\text{CE}}-y)^{\top}(\hat{p}_{\text{KD}}-p)$ remains negligible, keeping $\mathtt{CosSim}(\nabla_{z}\mathcal{L}_{\text{CE}},\nabla_{z}\mathcal{L}_{\text{KD}})$ close to zero (Figure 13-(right)). DHO thus maintains consistently positive gradient alignment, i.e., $\mathtt{CosSim}(\nabla_{\theta}\mathcal{L}_{\text{CE}},\nabla_{\theta}\mathcal{L}_{\text{KD}})>0$ (as shown in of Figure 4), enabling stable and conflict-free representation learning. See Algorithm 1 for the full training procedure of DHO.
Language-aware initialization and KD head alignment for VLM students.
@@
We initialize the parameters of the dual heads $h_{\text{CE}}$ and $h_{\text{KD}}$, i.e., $W_{\text{CE}},W_{\text{KD}},b_{\text{CE}},b_{\text{KD}}$, using Kaiming initialization [31]. In the case of VLM-to-VLM distillation, however, we can leverage the teacher’s text encoder $f_{\mathbb{T}}$. Following prior work on language-aware initialization [51], we initialize the weights as $W_{\text{CE}},W_{\text{KD}}\leftarrow[f_{\mathbb{T}}(t_{1}),\ldots,f_{\mathbb{T}}(t_{C})]^{\top}\in\mathbb{R}^{C\times d}$. Furthermore, we align the prediction logic of KD head with that of teacher as follows:
@@
where $w_{c}\in\mathbb{R}^{d}$ denotes the $c$-th row of $W_{\text{KD}}$. This alignment encourages the KD head to mimic the teacher’s cosine similarity-based prediction behavior, leading to more stable distillation and improved performance, as shown in Table 5.
Dual-heads interpolation for inference.
We propose a simple yet effective inference strategy for dual-head models by linearly interpolating:
$\hat{p}=\alpha\cdot\sigma(h_{\text{CE}}(z))+(1-\alpha)\cdot\sigma(h_{\text{KD}}(z)/\beta)\in\Delta^{C-1},$ (11)
where $\alpha\in[0,1]$ is an interpolation hyperparameter that balances the influence of the supervised and KD heads, and $\beta\in\mathbb{R}_{>0}$ is a temperature parameter that softens the KD logits. The final prediction is obtained via $\arg\max_{c\in[C]}\hat{p}_{c}$. This interpolation scheme is inspired by the mixture-of-experts (MoE) paradigm [41], where multiple predictive sources are combined, and their relative importance is adjusted based on task conditions. In our setting, $\alpha$ allows flexible weighting of the two heads depending on which source—supervised labels or teacher predictions—offers more reliable guidance for a given dataset. This balance can be tuned using a validation set to reflect dataset-specific supervision quality and teacher accuracy. See Algorithm 2 for the detailed inference process of DHO.
@@ 4 Experiments
We now elaborate the empirical effectiveness of the proposed method DHO. We first detail experimental setups in §4.1, and report the following empirical findings in §4.2:
  • [F1]: In comparison to single-head KD baselines, DHO is an effective framework for semi-supervised settings, with negligible additional computational cost.
  • [F2]: The improvement of DHO is attributed to gradient conflict mitigation, which leads to better feature representations compared to single-head KD baselines.
  • [F3]: The proposed dual-heads output probability interpolation for inference further improves the performance of DHO on top of its enhanced feature representations.
  • [F4]: DHO achieves state-of-the-art performance on ImageNet under low-shot semi-supervised settings, while using fewer parameters.
@@ 4.1 Experimental Setups
We use 11 diverse image classification datasets, including generic object recognition such as ImageNet [76], Caltech101 [26] and other widely-adopted benchmarks for fine-grained classification (Cars [48], Flowers102 [67], FGVCAircraft [59], OxfordPets [69]), domain-specific recognition (Food101 [6]), scene understanding (SUN397 [99]), texture analysis (DTD [17]), satellite imagery (EuroSAT [34]), and human actions (UCF101 [81]). See Appendix C for details.
@@
We first compare DHO with conventional single-head KD baselines; CE: we train a model only on labeled dataset $\mathcal{D}^{(l)}$ with cross entropy loss (Eq. 2), KD (logit): on unlabeled dataset $\mathcal{D}^{(u)}$ with logit distillation (Eq. 3), and KD (feature): on $\mathcal{D}^{(u)}$ with feature distillation [101]. We train a model on both $\mathcal{D}^{(l)}$ and $\mathcal{D}^{(u)}$ with CE+KD (logit) or CE+KD (feature): which combines CE with each KD variant with the balancing hyperparameter $\lambda$. We also consider existing dual-head KD approaches: SSKD [33] uses dual-heads, with each head trained on labeled and unlabeled sets respectively, assuming different data distributions between them, and DHKD [104] introduces binary KD loss working on logits before softmax to prevent neural collapse [68], but both methods predict labels using only $h_{\text{CE}}$. We further compare DHO with state-of-the-art methods on ImageNet with low-shot semi-supervised settings, including self and semi-supervised learning [118, 2, 13, 13, 3, 7, 7], CLIP-based-training [51, 55], co-training [74], KD [13], and zero-shot VLMs [112, 22].
Architecture choices for teacher and student models.
For zero-shot teachers, we use CLIP ResNet-50 [72] in few-shot settings and VIT-H/14 from DFN [22] in low-shot settings. For few-shot teachers, we adopt Tip-Adapter-F [115], a learnable adapter model, and denote the corresponding variant as DHO-F. Student models are initialized from pretrained checkpoints [87]. On ImageNet, to avoid label leakage, we either train ResNet-18 from scratch or use a self-supervised ResNet-50 from DINO [8]. For other datasets, we use ResNet-18 and MobileNetV2 [77] without such concerns. In low-shot settings, we use CLIP ViT-B/16 and ViT-L/14 [72]. See Table 6 for an overview.
@@
We consider few-shot (1, 2, 4, 8, 16-shot) and low-shot (1% or 10%) settings, treating the rest as unlabeled. Prompt templates follow Tip-Adapter [115]. As detailed in §3.2, we use linear heads for all methods; additional analysis on non-linear heads is provided in Section E.2. All methods use consistent hyperparameters, including $\zeta=0.01$, $\lambda=0.5$, and others. For KD (feature) baselines, we scale the MSE loss to match $\mathcal{L}_{\text{CE}}$ following [101]. $\alpha$ and $\beta$ of DHO are tuned on validation sets. When validation data is unavailable (e.g., ImageNet), we fix $\beta=0.5$ across all settings. We heuristically set $\alpha=0.4$ for zero-shot teachers and $\alpha=0.2$ for few-shot teachers, reflecting the latter’s higher reliability. In low-shot settings, we use $\alpha=0.5$ due to increased label availability. Results are from a single run due to limited compute. See Section B.2 for more details.
4.2 Experimental Results
Table 1: Results on ImageNet under few-shot semi-supervision using ResNet-18 and ResNet-50.
@@ 4.2 Experimental Results
DHO (zero-shot teacher [72]) - 51.8 52.4 52.6 53.3 54.5 - 61.0 62.1 62.5 63.7 64.7
DHO-F (few-shot teacher [115]) - 53.7 54.2 54.8 56.2 57.7 - 62.3 63.1 63.9 65.5 66.8
Teacher Models
CLIP [72] 50.0 - - - - - 60.3 - - - - -
Tip-Adapter-F [115] - 52.0 52.4 53.2 54.0 55.3 - 61.0 61.6 62.5 63.8 65.4
Figure 5: Results on 10 datasets under few-shot semi-supervision using ResNet-18 with zero-shot teacher [72].
Figure 6: Results on 10 datasets using ResNet-18 with either zero- or few-shot teacher [115].

### v2

@@ 2.2 Dual-Head Optimization (DHO)
Figure 3: The average cosine similarity and inner product over 10 datasets.
Gradient conflicts in SHO.
Logit distillation in SHO, as described in Section˜2.1, provides a simple approach for transferring knowledge from VLMs to task-specific models. However, we find that its performance gain is suboptimal, which we attribute to gradient conflicts between the supervised and KD loss signals which hinder effective feature learning (Yu et al., 2020a; Liu et al., 2021; Chen & Er, 2025). As illustrated in and of Fig.˜3, both the classifier weight $W$ of classifier $h$ and the parameter vector $\theta$ of feature extractor $g$ suffer from gradient conflicts: the cosine similarity between their respective gradients turns negative, i.e., $\mathtt{CosSim}(\nabla_{\theta}\mathcal{L}_{\text{CE}},\nabla_{\theta}\mathcal{L}_{\text{KD}})<0$ and $\mathtt{CosSim}(\nabla_{W}\mathcal{L}_{\text{CE}},\nabla_{W}\mathcal{L}_{\text{KD}})<0$, indicating misaligned optimization directions.
To understand this, we first analyze the gradient with respect to (w.r.t) the weight $W$ of a linear head $h$. The gradients w.r.t the classifier weight $W$ are $\nabla_{W}\mathcal{L}_{\text{CE}}=(\hat{p}-y)z^{\top}$ and $\nabla_{W}\mathcal{L}_{\text{KD}}=(\hat{p}-p)z^{\top}$, respectively. Their cosine similarity is proportional to $(\hat{p}-y)^{\top}(\hat{p}-p)\cdot\|z\|^{2}$:
$\mathtt{CosSim}(\nabla_{W}\mathcal{L}_{\text{CE}},\nabla_{W}\mathcal{L}_{\text{KD}})\propto(\hat{p}-y)^{\top}(\hat{p}-p)\cdot\|z\|^{2},$ (3)
which misaligns when $(\hat{p}-y)^{\top}(\hat{p}-p)<0$ due to prediction mismatch, falling below zero during training ( $\times$ in Fig.˜3), leading to gradient conflicts ( in Fig.˜3).
We observe that similar gradient conflicts happen on the parameters $\theta$ in the feature extractor $g$. Let the gradients w.r.t to the feature representation $z$ be $\nabla_{z}\mathcal{L}_{\text{CE}}=W^{\top}(\hat{p}-y)$ and $\nabla_{z}\mathcal{L}_{\text{KD}}=W^{\top}(\hat{p}-p)$, Then, applying the chain rule, the gradients with respect to the feature extractor parameters $\theta$ become $\nabla_{\theta}\mathcal{L}_{\text{CE}}=\nabla_{z}\mathcal{L}_{\text{CE}}\cdot\frac{\partial z}{\partial\theta}$ and $\nabla_{\theta}\mathcal{L}_{\text{KD}}=\nabla_{z}\mathcal{L}_{\text{KD}}\cdot\frac{\partial z}{\partial\theta}$. Similarly, the cosine similarity for the feature extractor parameters becomes:
$\mathtt{CosSim}(\nabla_{\theta}\mathcal{L}_{\text{CE}},\nabla_{\theta}\mathcal{L}_{\text{KD}})\propto(\hat{p}-y)^{\top}W\left(\frac{\partial z}{\partial\theta}\right)^{\top}\frac{\partial z}{\partial\theta}W^{\top}(\hat{p}-p)$ (4)
While the complexity of the Jacobian $\frac{\partial z}{\partial\theta}$ makes it difficult to theoretically guarantee gradient conflicts, we empirically observe that gradient conflicts occur in the parameters $\theta$ of the feature extractor $g$ ( in Fig.˜3), with $\mathtt{CosSim}(\nabla_{\theta}\mathcal{L}_{\text{CE}},\nabla_{\theta}\mathcal{L}_{\text{KD}})<0$ . Thus, we hypothesize that the gradient conflicts in the feature extractor $g$ are propagated from the classification head $h$.
@@
To mitigate this issue, we propose Dual Head Optimization (DHO) to decouple $\mathcal{L}_{\text{CE}}$ and $\mathcal{L}_{\text{KD}}$ via two independent prediction heads: $h_{\text{CE}}(z)=W_{\text{CE}}z+b_{\text{CE}}$ and $h_{\text{KD}}(z)=W_{\text{KD}}z+b_{\text{KD}}$, with $W_{\text{CE}},W_{\text{KD}}\in\mathbb{R}^{C\times d}$ and $b_{\text{CE}},b_{\text{KD}}\in\mathbb{R}^{C}$. The corresponding losses are:
@@
where the final loss combines both objectives as $\lambda\mathcal{L}_{\text{CE}}+(1-\lambda)\mathcal{L}_{\text{KD}}$.
Mitigation of gradient conflict in DHO.
In DHO, gradient conflicts in the classification head naturally disappear by decoupling the optimization of $W_{\text{CE}}$ and $W_{\text{KD}}$. This separation enables each head to learn distinct signals without interference, reducing prediction mismatch ( $\times$ of Fig.˜3). Let $\hat{p}_{\text{CE}}{=}\sigma(h_{\text{CE}}(z))$ and $\hat{p}_{\text{KD}}{=}\sigma(h_{\text{KD}}(z))$, then $\nabla_{z}\mathcal{L}_{\text{CE}}{=}W_{\text{CE}}^{\top}(\hat{p}_{\text{CE}}-y)$ and $\nabla_{z}\mathcal{L}_{\text{KD}}{=}W_{\text{KD}}^{\top}(\hat{p}_{\text{KD}}-p)$. The cosine similarity between gradients w.r.t $\theta$ in DHO is defined as:
@@
$\mathtt{CosSim}(\nabla_{\theta}\mathcal{L}_{\text{CE}},\nabla_{\theta}\mathcal{L}_{\text{KD}})\propto(\hat{p}_{\text{CE}}-y)^{\top}W_{\text{CE}}\left(\frac{\partial z}{\partial\theta}\right)^{\top}\frac{\partial z}{\partial\theta}W_{\text{KD}}^{\top}(\hat{p}_{\text{KD}}-p),$ (7)
where we empirically find that gradient conflicts in $\theta$ of the feature extractor $g$ for DHO are also resolved, i.e., maintaining positive gradient alignment throughout training (i.e., Eq.˜7 > 0; as shown in of Fig.˜3). It enables conflict-free representation learning, and leads to better feature representation compared to SHO, which empirically validated by linear evaluation in Section˜3.3. We defer an algorithm that describes the full training procedure of DHO to Alg.˜1.
2.3 Dual-head Interpolation
After training, we find that using only one of the two heads at inference is suboptimal (see Fig.˜6). Motivated by the mixture-of-experts paradigm (Jacobs et al., 1991), we adopt a simple yet effective inference rule for DHO that linearly interpolates the output probability vectors of two heads:
$\hat{p}_{\textbf{{DHO}}}=\alpha\,\sigma\!\big(h_{\text{CE}}(z)\big)+(1-\alpha)\,\sigma\!\big(h_{\text{KD}}(z)/\beta\big),$ (8)
where $\alpha\in[0,1]$ balances the supervised and KD heads, $\beta>0$ is a temperature that softens the KD logits, and the final prediction is $\arg\max_{c\in[C]}\hat{p}_{\textbf{{DHO}}}$. In practice, we tune $\alpha$ and $\beta$ on a validation set to reflect dataset-specific supervision quality and teacher accuracy, allowing the model to weight the more reliable source. See Alg.˜2 for the full inference procedure.
Effect of $\alpha$ and $\beta$.
We now demonstrate the effect of tuning $\alpha$ and $\beta$ using a validation set. Under mild assumptions, DHO $\varepsilon$-approximates SHO in $\ell_{1}$ by setting $\alpha=\lambda$ and $\beta=1$:
Assumption 1 ($\varepsilon$-convergence).
Assume that, after sufficient training, both heads converge to their respective target distributions with $\ell_{1}$-bounded error:
$\sup_{x}\|\sigma(h_{\mathrm{CE}}(z))-y\|_{1}\leq\varepsilon,\quad\sup_{x}\|\sigma\!(h_{\mathrm{KD}}(z))-p\|_{1}\leq\varepsilon,\quad\text{where}\quad\varepsilon\in\mathbb{R}_{>0}.$ (9)
Theorem 1 (Inference equivalence).
Under Assumption˜1, by setting $\alpha=\lambda$ and $\beta=1$, then $\|\hat{p}_{\textbf{{DHO}}}-\hat{p}_{\mathrm{SHO}}\|_{1}\leq\varepsilon,$ where $\hat{p}_{\mathrm{SHO}}$ is the output of SHO optimally trained with $\lambda$.
Details and proofs are deferred to Appendix˜B. Theorem˜1 implies that $\hat{p}_{\text{SHO}}$ trained with any $\lambda$ can be approximated by the dual-head interpolation in Eq.˜8, by setting $\alpha=\lambda$ and $\beta=1$. Here, $\lambda$ is a training hyperparameter of SHO, while $\alpha$ and $\beta$ are inference hyperparameters of DHO, allowing it to emulate SHO hyperparameter tuning without retraining.
Language-aware initialization for VLM students
@@ 2.3 Dual-head Interpolation
In the case of VLM-to-VLM distillation, we leverage the text encoder $f_{\mathcal{T}}$ of teachers when initializing the dual heads $h_{\text{CE}}$ and $h_{\text{KD}}$. Following prior work (Li et al., 2022), we initialize the weights as $W_{\text{CE}},W_{\text{KD}}\leftarrow[f_{\mathcal{T}}(t_{1}),\ldots,f_{\mathcal{T}}(t_{C})]^{\top}\in\mathbb{R}^{C\times d}$. We further align the prediction logic of KD head $h_{\text{KD}}$, with the cosine similarity-based approach of the teacher VLMs as follows:
@@
where $w_{c}\in\mathbb{R}^{d}$ denotes the $c$-th row of $W_{\text{KD}}$.
@@ 3.1 Experimental Setups
For in-distribution evaluation, we use ImageNet (Russakovsky et al., 2015) and 10 fine-grained datasets (Fei-Fei et al., 2004; Parkhi et al., 2012; Krause et al., 2013; Nilsback & Zisserman, 2008; Bossard et al., 2014; Maji et al., 2013; Xiao et al., 2010; Cimpoi et al., 2014; Helber et al., 2019; Soomro, 2012). To assess out-of-distribution (OOD) generalization, we use four ImageNet variants (Recht et al., 2019; Wang et al., 2019; Hendrycks et al., 2021a; b). See Appendix˜D for details.
@@
We compare DHO with conventional single-head KD baselines; CE: training only on labeled dataset $\mathcal{D}^{(l)}$ with cross entropy loss (Eq.˜1), KD (logit): on unlabeled dataset $\mathcal{D}^{(u)}$ with logit distillation (Eq.˜2), and KD (feature): on $\mathcal{D}^{(u)}$ with feature distillation (Yang et al., 2024a). We train on both $\mathcal{D}^{(l)}$ and $\mathcal{D}^{(u)}$ with CE+KD (logit) or CE+KD (feature): combining CE with each KD variant using balancing hyperparameter $\lambda$. We also consider dual-head KD approaches SSKD (He et al., 2021) and DHKD (Yang et al., 2024d), though for different purposes as detailed in Appendix˜A.
For in-distribution evaluation on ImageNet with low-shot settings, we compare against self and semi-supervised learning (Chen et al., 2020; Assran et al., 2021; 2022; Cai et al., 2022; Zheng et al., 2023), CLIP-based-training (Li et al., 2022; Liu et al., 2023b), co-training (Rothenberger & Diochnos, 2023), KD (Chen et al., 2020), and zero-shot VLMs (Zhai et al., 2023; Fang et al., 2023).
For OOD evaluation, we compare against VLM adaptation methods, including VPT (Jia et al., 2022), CoOp (Zhou et al., 2022b), PromptSRC (Khattak et al., 2023b), and CasPL (Wu et al., 2025).
@@
We evaluate across few-shot (1/2/4/8/16-shot) and low-shot (1%/10%) settings, treating remaining data as unlabeled. We adopt Tip-Adapter-F (Zhang et al., 2021) for few-shot teachers, denoting this variant as DHO-F. For a fair comparison, we use the same hyperparameters for DHO and SHO baselines (except $\alpha,\beta$); for other methods, we report the published results. When validation data is unavailable (e.g., ImageNet), we fix $\beta=0.5$ across all settings. We heuristically set $\alpha=0.4$ for zero-shot teachers and $\alpha=0.2$ for few-shot teachers, reflecting the latter’s higher reliability. In low-shot settings, we use $\alpha=0.5$ due to increased label availability. See Appendix˜C for details.
Table 1: Results on ImageNet under few-shot semi-supervision using ResNet-18 and ResNet-50. DHO consistently outperforms all baselines and even the teacher with ResNet-50 (e.g., +0.7/1.9/2.2/3.4/4.4% with a zero-shot teacher; +1.3/1.5/1.4/1.7/1.4% with a few-shot teacher).
@@
SSKD (He et al., 2021) - 42.5 46.2 48.0 50.6 52.0 - 55.2 58.1 60.0 62.3 64.0
DHKD (Yang et al., 2024d) - 19.7 23.5 23.4 23.7 26.8 - 25.6 34.8 42.7 49.2 55.2
DHO (Ours) - 51.8 52.4 52.6 53.3 54.5 - 61.0 62.1 62.5 63.7 64.7
DHO-F (Ours) - 53.7 54.2 54.8 56.2 57.7 - 62.3 63.1 63.9 65.5 66.8
Teacher Models (Resnet-50)
CLIP 60.3 - - - - - 60.3 - - - - -
Tip-Adapter-F - 61.0 61.6 62.5 63.8 65.4 - 61.0 61.6 62.5 63.8 65.4
Figure 4: Results on 10 datasets under few-shot semi-supervision using ResNet-18 with zero-shot teacher.
Figure 5: Results on 10 datasets using ResNet-18 with either zero- or few-shot teacher.
Table 2: Results on ImageNet and its distribution-shifted variants. $\dagger$: backbone frozen during training. DHO consistently improves VLM adaptation methods and achieves SoTA results on OOD benchmarks.
Method Labeled Data Teacher Model Unlabeled Data Val V2 Sketch R A Avg
CLIP Zero-shot
CLIP - - - 66.7 60.8 46.2 74.0 47.8 59.1
Linear Evaluation
Linear Evaluation 1% - - 72.8 64.2 47.0 74.6 48.6 61.4
DHO 1% - 73.4 65.3 48.3 75.8 49.4 62.4
DHO 1% 74.6 66.4 49.2 76.1 49.8 63.2
Visual Prompt Tuning
VPT 1.25% - - 73.6 64.6 47.7 75.2 48.7 62.0
VPT+DHO 1.25% - 73.6 65.3 48.7 75.9 49.1 62.5
VPT+DHO 1.25% 75.1 66.8 50.0 76.9 50.5 63.9
VLM Text-encoder Prompt Tuning
CoOp 1.25% - - 71.5 64.2 48.0 75.2 49.7 61.7
CoOp+CasPL 1.25% 71.9 64.3 48.3 76.0 - -
CoOp+DHO 1.25% - 72.8 65.5 49.3 76.4 49.5 62.7
CoOp+DHO 1.25% 73.4 66.2 49.5 77.0 50.5 63.3
VLM Multimodal Prompt Tuning
PromptSRC 1.25% - - 71.3 64.4 49.6 77.8 50.9 62.8
PromptSRC+CasPL 1.25% 72.8 65.7 49.7 77.9 - -
PromptSRC+DHO 1.25% - 73.0 65.3 49.5 77.8 51.3 63.4
PromptSRC+DHO 1.25% 73.6 66.1 49.8 78.1 51.0 63.7
3.2 Main Results
Effectiveness of DHO compared to conventional KD baselines.
Tab.˜1 presents results on ImageNet under few-shot semi-supervision using ResNet-18/ResNet-50 student models and ResNet-50 VLM teachers. DHO consistently outperforms all baselines across all settings. Other dual-head methods (i.e., SSKD/DHKD), not designed for few-shot semi-supervision, even underperform CE+KD on logits. Notably, with ResNet-50, DHO outperforms the teacher in every few-shot setting (e.g., +0.7/1.8/2.2/3.4/4.4% or +1.3/1.5/1.4/1.7/1.4% with zero-shot or few-shot teachers).
Next, we evaluate DHO on 10 additional datasets using the ResNet-18 student and ResNet-50 VLM teachers. In Fig.˜4, we observe that DHO also consistently outperforms all baseline methods across 10 datasets, while relative rankings between baselines vary across datasets. Fig.˜5 further demonstrates that the few-shot teacher is more effective than using a zero-shot teacher for DHO. Remarkably, the ResNet-18 student model trained with DHO achieves better performance than the ResNet-50 teacher model in most cases, demonstrating knowledge transfer capability of DHO.
Method Architecture Params 1% 10%
(M) (%) (%)
Self and Semi-supervised Learning
MSN ViT-B/4 86 75.7 80.2
Semi-ViT ViT-L/14 307 77.3 83.3
Semi-ViT ViT-H/14 632 80.0 84.3
CLIP-based Training
CLIP ViT-B/16 86 74.3 80.4
REACT ViT-B/16 86 76.1 80.8
REACT (Gated-Image) ViT-B/16 129 77.4 81.8
CLIP ViT-L/14 304 80.5 84.7
REACT ViT-L/14 304 81.6 85.1
REACT (Gated-Image) ViT-L/14 380 81.6 85.0
Co-training based Methods
CT Multi-arch (2) 608 80.1 85.1
MCT Multi-arch (2) 608 80.7 85.2
CT Multi-arch (4) 1071 80.0 84.8
MCT Multi-arch (4) 1071 80.5 85.8
Knowledge Distillation
SimCLR v2 distill ResNet-50 (2×+SK) 140 75.9 80.2
SimCLR v2 self-distill ResNet-154 (3×+SK) 795 76.6 80.9
CE + KD (logit) ViT-B/16 86 79.8 80.4
CE + KD (logit) ViT-L/14 304 83.1 83.6
DHO ViT-B/16 86 81.6 82.8
DHO ViT-L/14 304 84.6 85.9
Zero-shot VLMs
SigLIP ViT-SO400M/14 400 83.1 -
DFN ViT-H/14 632 84.4 -
DHO achieves SoTA performance on ImageNet under low-shot settings.
We compare DHO to previous state-of-the-art (SoTA) methods on ImageNet under 1% and 10% labeled data. All results are taken from published papers, except for DHO and CE+KD (logit). As shown in Tab.˜3, DHO with ViT-L/14 surpasses the previous SoTA (e.g., +3.0%/+0.1%), while using fewer parameters (e.g., 76M/767M), on both 1%/10% labeled data. Notably, CE + KD (logit) outperforms semi-supervised methods with the same parameters, demonstrating that they are suboptimal compared to methods leveraging the rich representations of VLMs.
DHO achieves SoTA performance on ImageNet OOD benchmarks.
We evaluate DHO across various VLM adaptation approaches. As shown in Tab.˜2, DHO consistently improves different adaptation methods. Compared to CasPL, DHO exhibits superior performance across both in-distribution (Val) and out-of-distribution (V2/Sketch/R/A) benchmarks, establishing new SoTA OOD results in semi-supervised setting. This suggests that joint training with labeled supervision and teacher distillation provides a more effective strategy than the sequential approach of CasPL, aligning with our hypothesis on DHO for resolving gradient conflicts. See Section˜E.5 for results with fully trained models.
Addtional results.
See Appendix˜E for additional results of MobileNetV2, or comparison to PCGrad (Yu et al., 2020b) and category-aware KD methods (Zhao et al., 2022; Lv et al., 2024).
3.3 Analysis